OpenStudy (anonymous):

Line R is represented by the following equation: x + y = 2 Which equation completes the system that is satisfied by the solution (1, 1)? 2x + y = 2 4x – 2y = 2 2x – 2y = 2 x + y = 4

OpenStudy (xapproachesinfinity):

all you need to do is plug in the x=1 and y=1 for all of them and see which words example for the first 2(1)+(1)=2+1=3 but three is not 2 so does not work continue like

OpenStudy (anonymous):

that does not help me at all i don't get what you said